Air at 50% relative humidity is cooled isobarically at 1 atm absolute from 90⁰C to 25⁰C.
a) Estimate the dew point and degrees of superheat of the air at 90⁰C.
b) How much water condenses (mol) per cubic meter of feed gas?
c) Suppose a sample of the 90⁰C air is put in a closed variable-volume chamber
containing a mirror and the pressure is raised at constant temperature until a mist forms on the
mirror. At what pressure (atm) would the mist form? (Assume ideal gas behavior.
Answer #1

Here is the explanation-

we know th Relative hum adi h, ,,Napor pressure ōfwater at 70 c. is s-25. 6mmH7 50% , nd Napor pressure of water at 1o cis 76 0.O mm Nou For (a) h-50 % 2 0.50 x S2구,76 mm Ha 760.0 mm Ha mol -0. 346 Hao/rno, Deuo point Tdp 20 P= PXTdP - 0.346 x 140.0=262.9 mm Hg Im-the table, the temperature at 26 2.9 mmHg is 72.7 с . So , the dew point is 72-7 c Degree o superheat 90-72. 1.3°c of superheat for (b) Basis: 23 33-6 mo m3 363k2.4 L (STP) Pano (25%) 23.756 Saturation 0.0313 mol H2O/mol Condition : Dry air bal.6s4 (334) n (1-0.0313) Total mol balance= 33.6-22-7 tn, 7 60 m22. mol n2= 10.9 mo, H20 condense/m 3 For () 0.346 YH20 2, 00 atm Hence, at 2.o0 atm the mist fll form.